# Inflationary attractors and radiative corrections in light of ACT

### v1

@@ Abstract
In light of the recent results from the Atacama Cosmology Telescope (ACT), which have provided a notable shift in the constraints on $(n_{s},r)$ and placed several otherwise viable models of inflation in tension with the latest data, we investigate the possible effects that radiative corrections can have on $\xi$-attractor and $\alpha$-attractor models of inflation. These models, which share much in common with Starobinsky-Higgs inflation, have likewise been put under pressure by these results. We find that percent (and even sub-percent) level radiative corrections can easily shift both of these classes of inflation models comfortably into the regions of parameter space favoured by the most recent constraints. However, the flexibility under such corrections (which one can argue should be generally expected) calls into question to what extent it is possible to precisely pin down model-specific predictions for important cosmological observables.
@@ I Introduction
Cosmic inflation has widely been considered to be the best paradigm to model the early universe ever since its introduction [1, 2, 3, 4, 5, 6, 7, 8]. In this recent era of precision cosmology, we have accurately measured the statistical features of the primordial density fluctuations that seed cosmic structure and found that they are in remarkable agreement with generic inflationary predictions [9, 10, 11, 12]. While there are a plethora of modeling constructs [13], the Starobinsky-Higgs model of inflation [1, 14, 15, 16, 17, 18, 19, 20, 21] has been accorded special significance as it can be viewed as simply an instantiation of GR with an $R^{2}$ correction at high energies, or as the standard model Higgs with a non-minimal coupling to gravity [13]. Furthermore, this model’s predictions have been shown to coincide with a number of other models [22, 23] and have consistently shown excellent compatibility with Cosmic Microwave Background (CMB) data [9].
Given these prevailing attitudes, the latest results from the Atacama Cosmology Telescope (ACT) [24], in combination with data from Planck [25], BICEP/Keck [26], and DESI [27], have generated a notable amount of discussion as this model now lies almost entirely outside the posterior constraints on $(n_{s},r)$. Recent work includes exploring the implications that radiative corrections, reheating, or other higher order effects might have on this model [28, 29, 30, 31, 32, 33, 34, 35], as well as other inflation constructs [36, 37, 38, 39, 40, 41, 42, 43, 44]. Here, we explore the impact that radiative corrections can have on various inflationary attractors known as $\xi$-attractors [22] and $\alpha$-attractors [23], both of which can be viewed as generalizations of Starobinsky-Higgs model as they have similar predictions and contain the Higgs-Starobinsky model as a subset for certain parameter choices. Similar to [28] which studied the impact of radiative corrections to Higgs-like inflation, we find that small, percent (and even sub-percent) level radiative correction can induce significant changes in observable predictions for these broader classes of models, and easily bring them into agreement with the latest constraints.
The paper proceeds as follows. Sec. (II) discusses the $\xi$-attractor and $\alpha$-attractor models. Sec. (III) follows [28, 45] in introducing a UV model agnostic way of phenomenologically parameterizing radiative corrections that might result from one loop corrections arising from the full spectrum of particle content at high energies and demonstrates that relatively small corrections can have a significant impact on observables. Sec. (IV) concludes with a discussion of these results and their implications for assessing inflation.
@@ II.1 $\xi$-attractors
The case of $\Omega(\varphi)=1+\xi\varphi^{2}$ and $V(\varphi)=\lambda\varphi^{4}$ (i.e., $f(\varphi)=\varphi^{2}$) corresponds to Starobinksy/Higgs-like inflation [14], but this class of models is far more general and can provide viable inflationary models for many choices of $f(\varphi)$. Indeed, it can be thought of as a generalization to the historically important chaotic inflation [8] models to include a non-minimal coupling to gravity [22].
@@ II.2 $\alpha$-attractors
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}R-K(\varphi)(\partial\varphi)^{2}-V(\varphi)\right].$ (6)
Depending on the choice of $K(\varphi)$, finding the canonically normalize field $\chi$ leads to the following analytic expressions for the potentials,
@@
where Eq. (7) represents so-called “T-models” and Eq. (8) represents so-called “E-models”. In the case of E-models, the choices $p=2$ and $\alpha=1$ correspond to the Starobinsky model. We set $p=2$ for the remainder of this paper. While the potentials are formally similar to the Starobinsky potential expressed in the Einstein frame, note here that there is no need to actually do the conformal transformation as we did with the $\xi$-attractors given that the field is not ever directly coupled to $R$ and/or never includes a higher order $R^{2}$ term as their origin in the simplest incarnations can be understood to come from modifications to the kinetic term (see the discussion in [11, Sec. 10]). The only necessary step is to canonically normalize the field.
@@ II.3 Baseline observable predictions
During inflation, quantum mechanical variations in the value of the field driving inflation will produce a spectrum of scalar and tensor perturbations in the early universe. The amplitude of scalar perturbations has been measured precisely to be $A_{\mathrm{s}}\simeq 2.1\times 10^{-9}$[9]. The primary observable predictions used to constrain inflation models are the tensor-to-scalar ratio $r\equiv A_{\mathrm{t}}/A_{\mathrm{s}}$, which quantifies the ratio of the amplitudes of scalar and the (yet to be detected) tensor perturbations, and the scalar spectral index $n_{s}\left(k_{*}\right)\equiv 1+\left.\left(d\ln\mathcal{P}_{\mathcal{R}}/d\ln k\right)\right|_{k_{*}}$, which quantifies the scale dependence of the scalar power spectrum $\mathcal{P}_{\mathcal{R}}$ at the pivot scale $k_{*}$. During standard scenarios where inflation is driven by the scalar field potential, the universe enters a quasi-de Sitter expansion that is well-described by the slow-roll approximation. This slow-roll approximation is valid when $\epsilon,\eta\ll 1$, where
$\epsilon\equiv\frac{1}{2}\left(\frac{1}{U(\chi)}\frac{\mathrm{d}U(\chi)}{\mathrm{d}\chi}\right)^{2},\quad\eta\equiv\frac{1}{U(\chi)}\frac{\mathrm{d}^{2}U(\chi)}{\mathrm{d}\chi^{2}}.$ (9)
Inflation ends when $\epsilon\simeq 1$. This allows us to calculate the number of e-folds $N(\chi)$ between the end of inflation and some earlier initial time when the fluctuations exit the horizon and freeze in the CMB,
$N(\chi)=\log\left(\frac{a_{\mathrm{e}}}{a_{\mathrm{cmb}}}\right)=\int_{\chi_{\mathrm{e}}}^{\chi_{\mathrm{cmb}}}\frac{d\chi}{\sqrt{2\epsilon(\chi)}}.$ (10)
Inflation generally needs $N\simeq 55\pm 5$ e-folds to provide a viable description of the early universe. Crucially, the observable quantities $r$ and $n_{s}$ can then be directly calculated from the slow-roll parameters, leading to
$n_{s}-1\simeq-6\epsilon\left(\chi_{\mathrm{cmb}}\right)+2\eta\left(\chi_{\mathrm{cmb}}\right),\quad r\simeq 16\epsilon\left(\chi_{\mathrm{cmb}}\right),$ (11)
@@
The $\xi$-attractors and $\alpha$-attractors have well-known predictions for $(n_{s},r)$. At $\xi\ll 1$, the predictions for the $\xi$-attractors reduce to those of their minimally coupled cousins. This part of the parameter space is highly disfavored as minimally coupled polynomial chaotic models of inflation generally predict $r$ much higher than current constraints. This is not in and of itself surprising. Inflation belongs to the realm of beyond the standard model particle physics and it would not be at all surprising if, e.g., a non-minimal coupling was needed to adequately account for observations as these are generically expected to be generated from radiative loop corrections [51]. [^2] Interestingly, while dark energy occurs in the IR (rather than UV), recent constraints from the late time universe are likewise beginning to favour more complicated/non-minimally coupled scalar field models given that simpler constructs like minimally coupled quintessence cannot describe current constraints on the expansion history as well [52, 53, 54, 55, 56, 57]. Around $\xi\simeq\mathcal{O}(1)$ they all rapidly converge to an attractor point in $(n_{s},r)$ which coincides with Starobinsky inflation,
$n_{s}=1-\frac{2}{N},\quad r=\frac{12}{N^{2}}.$ (12)
@@
As we can see, $N\simeq 55$ leads to $n_{s}\simeq 0.964$, which now lies outside the posterior constraints with a mean value of $n_{s}\simeq 0.974$.
@@ III.1 Modeling quantum corrections
Figure 1: Starobinsky/Higgs potential (solid blue line) compared with loop corrected effective potentials where the loop correction has been introduced to the non-minimally coupled quartic potential in the Jordan frame (dashed green line) and then transformed to the Einstein frame and the Starobinsky/E-model $\alpha$-attractor potential which is already defined in the Einstein frame (dashed yellow line). This produces a small, but noticable difference in the resulting effective potential.
@@
$\lambda(\varphi)=\lambda\left(\mu_{0}\right)+\sum_{k=1}^{\infty}\frac{\beta_{k}}{k!}\log^{k}\left(\frac{\varphi}{\mu_{0}}\right),$ (14)
where $\mu_{0}$ is the renormalization scale and $\beta_{k}$ is $k$th derivative of the $\beta$-function at this scale. Of course, the exact expressions for these parameters depend on knowing the exact particle content of the UV theory, but generally speaking Eq. (14) is a model independent way of characterizing radiative corrections as this functional form naturally emerge from evaluating one-loop diagrams [58, 45, 28, 59, 13].
Following [45, 28], we work in a model agnostic manner and consider the general phenomenology that can emerge from loop corrections to the inflation models selected here. That is, we take the first ($k=1$) term in the expansion from Eq. (14) as the dominant correction, leading to an effective $\lambda(\varphi)$,
@@
The calculation strategy of [45, 28] is to consider the one-loop effective potential $V(\varphi)$ by incorporating the running $\lambda(\varphi)$, and then to transform this new effective potential from the Jordan frame to the Einstein frame and canonically normalize the field to calculate inflationary observables (see [45, fn. 2]). We follow this same strategy for computing quantum corrections to the $\xi$-attractor models, resulting in an effective potential
@@
which can then be inserted into Eqs. (7) and (8) in the same manner just described. We follow this computational strategy for both $\alpha$-attractor models (see [60] for an example discussion of radiative corrections in $\alpha$-attractors). It is important to note that this will lead to different effective potentials for $\xi$-attractor and $\alpha$-attractor models whose uncorrected potentials coincide. For example, Fig. 1 depicts the Starobinsky model (which is identical to an $\alpha$-attractor E-model with $p=2$ and $\alpha=1$) and a Higgs-like model with a quartic Jordan frame potential non-minimally coupled to gravity ($f(\varphi)=\varphi^{2}$ in the $\xi$-attractor language). Transforming the uncorrected Higgs-like model into the Einstein frame results in an essentially identical potential to the uncorrected E-model/Starobinsky model. However, implementing a $\delta=0.025$ correction to the Higgs-like potential and then transforming to the Einstein frame results in a different effective potential than implementing a $\delta=0.025$ correction to the E-model potential $U(\chi)$. This is because the loop correction will be affected by the conformal transformation in the Higgs-like $\xi$-attractor, whereas the loop corrections will not undergo any transformation with the $\alpha$-attractors as they already originate in the Einstein frame, despite the formal resemblances with other similar models that are conceived of as non-minimally coupled to gravity. [^3] See [61, 62, 49, 63, 64] for related discussion concerning the classical equivalence and possible quantum non-equivalence of the Jordan and Einstein frames.

### v3

@@ Abstract
In light of the recent results from the Atacama Cosmology Telescope (ACT), which have provided a notable shift in the constraints on $(n_{s},r)$ and placed several otherwise viable models of inflation in tension with the latest data, we investigate the possible effects that radiative corrections can have on $\xi$-attractor and $\alpha$-attractor models of inflation. These models, which share much in common with Starobinsky inflation, have likewise been put under pressure by these results. We find that percent (and even sub-percent) level radiative corrections can easily shift both of these classes of inflation models comfortably into the regions of parameter space favoured by the most recent constraints. However, the flexibility under such corrections calls into question to what extent it is possible to precisely pin down model-specific predictions for important cosmological observables.
@@ I Introduction
Cosmic inflation has widely been considered to be the best paradigm to model the early universe ever since its introduction [1, 2, 3, 4, 5, 6, 7, 8]. In this recent era of precision cosmology, we have accurately measured the statistical features of the primordial density fluctuations that seed cosmic structure and found that they are in remarkable agreement with generic inflationary predictions [9, 10, 11, 12]. While there are a plethora of modeling constructs [13], the Starobinsky [1, 14, 15] and Higgs-like models of inflation [16, 17, 18, 19, 20] have been accorded special significance as one can be viewed as simply an instantiation of GR with an $R^{2}$ correction at high energies, while the other can be viewed as the standard model Higgs with a non-minimal coupling to gravity. Furthermore, these models’ predictions have been shown to coincide with a number of other models of interest [21, 22, 23, 24] and have consistently shown excellent compatibility with Cosmic Microwave Background (CMB) data.
Given these prevailing attitudes, the latest results from the Atacama Cosmology Telescope (ACT) [25], in combination with data from Planck [26], BICEP/Keck [27], and DESI [28], have generated a notable amount of discussion as this model now lies almost entirely outside the posterior constraints on $(n_{s},r)$. Recent work includes exploring the implications that radiative corrections, reheating, or other higher order effects might have on the Starobinsky or Higgs-like models [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], as well as other inflation constructs [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. Here, we explore the impact that radiative corrections can have on various inflationary attractors known as $\xi$-attractors [21] and $\alpha$-attractors [22], the first of which can be viewed as a generalization of the Higgs-like model and the second of which can be viewed as a generalization of the Starobinsky model (for more on these attractors and their relation to each other see [67, 68]). Similar to [29] which studied the impact of radiative corrections to Higgs-like inflation, we find that small, percent (and even sub-percent) level radiative correction can induce significant changes in observable predictions for these broader classes of models, and easily bring them into agreement with the latest constraints.
The paper proceeds as follows. Sec. (II) discusses the $\xi$-attractor and $\alpha$-attractor models. Sec. (III) follows [29, 69, 70] in introducing a UV model agnostic way of phenomenologically parameterizing radiative corrections that could result from one loop corrections arising from the full spectrum of particle content at high energies and demonstrates that relatively small corrections can have an impact on observables. Sec. (IV) concludes with a discussion of these results and their implications for assessing inflation.
@@ II.1 $\xi$-attractors
The case of $\Omega(\varphi)=1+\xi\varphi^{2}$ and $V(\varphi)=\lambda\varphi^{4}$ (i.e. $f(\varphi)=\varphi^{2}$) corresponds to Higgs-like inflation [16], but this class of models is far more general and can provide viable inflationary models for many choices of $f(\varphi)$. Indeed, it can be thought of as a generalization to the historically important chaotic inflation [8] models to include a non-minimal coupling to gravity [21].
@@ II.2 $\alpha$-attractors
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}R-\frac{1}{2}K(\varphi)(\partial\varphi)^{2}-V(\varphi)\right].$ (6)
Two well-known $\alpha$-attractor models are so-called “T-models” and “E-models”, where T-models correspond to $K(\varphi)=1/(1-\varphi^{2}/6\alpha)^{2}$ and E-models correspond to $K(\varphi)=3\alpha/2\varphi^{2}$ [11].
These can be cast into the more familiar canonical form
by finding the canonically normalize field $\chi$. This leads to the following analytic expressions for the potentials,
@@
where Eq. (8) represents the T-models and Eq. (9) represents the E-models. In the case of E-models, the choices $p=2$ and $\alpha=1$ correspond to the Starobinsky model. We set $p=2$ for the remainder of this paper. While the potentials are formally similar to the Starobinsky potential expressed in the Einstein frame, note here that there is no need to actually do the conformal transformation as we did with the $\xi$-attractors given that the field is not ever directly coupled to $R$ and/or never includes a higher order $R^{2}$ term as their origin in the simplest incarnations can be understood to come from modifications to the kinetic term (see the discussion in [11, Sec. 10]). The only necessary step is to canonically normalize the field.
@@ II.3 Baseline observable predictions
During inflation, quantum mechanical variations in the value of the field driving inflation will produce a spectrum of scalar and tensor perturbations in the early universe. The amplitude of scalar perturbations has been measured precisely to be $A_{\mathrm{s}}\simeq 2.1\times 10^{-9}$ at the comoving CMB pivot scale $k_{\ast}=0.05\,\mathrm{Mpc}^{-1}$ [9]. The primary observable predictions used to constrain inflation models are the tensor-to-scalar ratio $r\equiv A_{\mathrm{t}}/A_{\mathrm{s}}$, which quantifies the ratio of the amplitudes of scalar and the (yet to be detected) tensor perturbations, and the scalar spectral index $n_{s}\left(k_{*}\right)\equiv 1+\left.\left(d\ln\mathcal{P}_{\mathcal{R}}/d\ln k\right)\right|_{k_{*}}$, which quantifies the scale dependence of the scalar power spectrum $\mathcal{P}_{\mathcal{R}}$. Finally, there is also a third observable that has generated some recent attention; the running of the spectral index $\alpha_{s}\equiv dn_{s}/d\ln k\big|_{k_{*}}$. During standard scenarios where inflation is driven by the scalar field potential, the universe enters a quasi-de Sitter expansion that is well-described by the slow-roll approximation. This slow-roll approximation [76, 77] is valid when $\epsilon_{\mathrm{v}},\eta_{\mathrm{v}},\xi_{\mathrm{v}}^{2}\ll 1$, where
$\epsilon_{\mathrm{v}}$ $\equiv\frac{1}{2}\!\left(\frac{1}{U(\chi)}\frac{\mathrm{d}U(\chi)}{\mathrm{d}\chi}\right)^{\!2},\quad\eta_{\mathrm{v}}\equiv\frac{1}{U(\chi)}\frac{\mathrm{d}^{2}U(\chi)}{\mathrm{d}\chi^{2}},$ (10)
$\xi_{\mathrm{v}}^{2}$ $\equiv\frac{1}{U(\chi)^{2}}\,\frac{\mathrm{d}U(\chi)}{\mathrm{d}\chi}\,\frac{\mathrm{d}^{3}U(\chi)}{\mathrm{d}\chi^{3}}.$
Inflation ends when $\epsilon_{\mathrm{v}}\simeq 1$. This allows us to calculate the number of e-folds $N(\chi)$ between the end of inflation and some earlier initial time when the fluctuations exit the horizon and freeze in the CMB,
$N(\chi)=\log\left(\frac{a_{\mathrm{e}}}{a_{\mathrm{cmb}}}\right)=\int_{\chi_{\mathrm{e}}}^{\chi_{\mathrm{cmb}}}\frac{d\chi}{\sqrt{2\epsilon(\chi)}},$ (11)
where $a_{\mathrm{e}}$ is the scale factor of the universe when inflation ends and $a_{\mathrm{cmb}}$} is the scale factor when modes exit the horizon. Inflation generally needs $N\simeq 55\pm 5$ e-folds to provide a viable description of the early universe. Crucially, the observable quantities $r$, $n_{s}$, and $\alpha_{s}$ can then be directly calculated from the slow-roll parameters, leading to
[…]
@@
The $\xi$-attractors and $\alpha$-attractors have well-known predictions for $(n_{s},r)$. At $\xi\ll 1$, the predictions for the $\xi$-attractors reduce to those of their minimally coupled cousins. This part of the parameter space is highly disfavored as minimally coupled polynomial chaotic models of inflation generally predict $r$ much higher than current constraints. This is not in and of itself surprising. Inflation belongs to the realm of beyond the standard model particle physics and it would not be at all surprising if e.g. a non-minimal coupling was needed to adequately account for observations as these are generically expected to be generated from radiative loop corrections [78]. [^2] Interestingly, while dark energy occurs in the IR (rather than UV), recent constraints from the late time universe are likewise beginning to hint at more intricate scalar field models that feature non-minimal interactions as the simplest quintessence models do not describe current constraints on the expansion history quite as well [79, 80, 81, 82, 83, 84]. Of course, any non-trivial interactions will also lead to other ancillary gravitational consequences, which may bring such models into tension with other datasets. Regardless, if one attempts to model either the early or late time universe with a single scalar field, various datasets are pointing towards more intricate scalar field models. Around $\xi\simeq\mathcal{O}(1)$ they all rapidly converge to an attractor point in $(n_{s},r)$ which coincides with Starobinsky inflation,
$n_{s}\simeq 1-\frac{2}{N},\quad r\simeq\frac{12}{N^{2}}.$ (13)
@@
As we can see, $N\simeq 55$ leads to $n_{s}\simeq 0.964$, which now lies outside the posterior constraints with a mean value of $n_{s}=0.974\pm 0.003$, indicating a $\simeq 3\sigma$ tension. [^3] The exact number of e-folds varies depending on the model considered and the assumptions around reheating. $N\simeq 55\pm 5$ represents a shorthand, model agnostic approximation that generally captures many of the most common models of inflation [85]. Even given the reheating analysis of [86, 87] which gives $N$ closer to the upper range of $N\simeq 60$ for Higgs inflation still results in an approximately $\simeq 2\sigma$ tension.
@@ III.1 Modeling quantum corrections
Figure 1: Higgs-like (with $\xi=10$) potential (solid blue line) compared with loop corrected effective potentials for the Higgs-like potential (dashed green line) and E-model $\alpha$-attractor (with $\alpha=1$) potential (dashed yellow line). While the Higgs-like potential and E-model potential at the chosen values for $\xi$ and $\alpha$ will coincide with the Starobinsky potential at to a very high degree of approximation at tree level, the radiatively corrected versions of these potentials differ more notably because in the Higgs-like case one must performs the conformal transformation after finding the effective potential whereas with the E-models one does not perform a conformal transformation.
@@
$\lambda(\varphi,\mu_{0})=\lambda\left(\mu_{0}\right)+\sum_{k=1}^{\infty}\frac{\beta_{k-1}}{k!}\log^{k}\left(\frac{\varphi}{\mu_{0}}\right),$ (15)
where $\mu_{0}$ is the renormalization scale and $\beta_{k}$ is $k$th derivative of the $\beta$-function at this scale. Generally speaking, Eq. (15) is a model independent way of characterizing radiative corrections as this functional form naturally emerge from evaluating one-loop diagrams [88, 69, 29, 89, 13, 73]. However, the exact expressions for these parameters depend on knowing the exact particle content of the UV theory, and indeed, depending on the assumptions on makes regarding UV physics, such corrections to inflationary observables could either be negligible or significant.
To just briefly consider an example, as discussed in [90, 91, 92, 73], there are numerous possible prescriptions for choosing the renormalization scale $\mu_{0}$. To say a bit more, $\mu$ is not truly constant and is more properly considered a function of the field $\mu(\varphi)$. For example, with one of the prescription (“prescription I”) that has been considered (chosen for its simplicity in the Einstein frame), $\mu(\varphi)$ evolves the same way as the masses of the particles being considered and merely changes the normalization of the potential without affecting its shape, meaning that inflationary observables will not change. On the other hand, another popular prescription (“prescription II”) motivated by quantization in the Jordan frame sees $\mu(\varphi)$ changing significantly during inflation; consequently, the shape of the potential is modified and radiative corrections can be significant to observable quantities (see the excellent discussion in [92] for more details). When one chooses the first prescription (along with a few other assumptions about the theory’s UV completion), it has been shown that radiative are negligible for Higgs inflation, $\xi$-attractors, and $\alpha$-attractors [90]; yet, as also discussed there, these results do not apply when the second renormalization prescription is chosen as these choices correspond to different UV completions and can lead to observable consequences for inflation observables [92, 91, 73]. Given our almost complete lack of knowledge concerning physical processes at these energy scales and the unsettled theoretical debates concerning UV completion, here we will not adopt an explicit UV completion and will pursue a different strategy.
Following [69, 29, 93, 70], we work in a model agnostic manner and consider the general phenomenology that can emerge from radiative corrections to the inflation models selected here. That is, we take the first ($k=1$) term in the expansion from Eq. (15) as the dominant correction, leading to an effective $\lambda(\varphi)$,
@@
The calculation strategy of [69, 29] is to consider the one-loop effective potential $V(\varphi)$ by incorporating the running $\lambda(\varphi)$, and then to transform this new effective potential from the Jordan frame to the Einstein frame and canonically normalize the field to calculate inflationary observables (see [69, fn. 2]). We follow this same strategy for computing quantum corrections to the $\xi$-attractor models, resulting in an effective potential
@@
which can then be inserted into Eqs. (8) and (9) in the same manner just described. We follow this computational strategy for both $\alpha$-attractor models (see [94] for an example discussion of radiative corrections in $\alpha$-attractors). It is important to note that this will lead to different effective potentials for $\xi$-attractor and $\alpha$-attractor models whose uncorrected potentials coincide. For example, Fig. 1 depicts the Starobinsky model (which is identical to an $\alpha$-attractor E-model with $p=2$ and $\alpha=1$) and a Higgs-like model with a quartic Jordan frame potential non-minimally coupled to gravity ($f(\varphi)=\varphi^{2}$ in the $\xi$-attractor language). Transforming the uncorrected Higgs-like model into the Einstein frame results in an essentially identical potential to the uncorrected E-model/Starobinsky model. However, implementing a $\delta=0.02$ correction to the Higgs-like potential and then transforming to the Einstein frame results in a different effective potential than implementing a $\delta=0.02$ correction to the E-model potential $U(\chi)$. This is because the loop correction will be affected by the conformal transformation in the Higgs-like $\xi$-attractor, whereas the loop corrections will not undergo any transformation with the $\alpha$-attractors as they already originate in the Einstein frame, despite the formal resemblances with other similar models that are conceived of as non-minimally coupled to gravity. [^4] See [95, 96, 74, 97, 98] for related discussion concerning the classical equivalence and possible quantum non-equivalence of the Jordan and Einstein frames.